### v5

# S2ST-Omni: An Efficient Multilingual Speech-to-Speech Translation Framework via Seamless Speech-Text Alignment and Progressive Fine-tuning

Abstract
Despite recent advances in multilingual speech-to-speech translation (S2ST), several critical challenges persist: 1) achieving high-quality translation remains a major hurdle, and 2) most existing methods heavily rely on large-scale parallel speech corpora, which are costly and difficult to obtain. To address these issues, we propose S2ST-Omni, an efficient and scalable framework for multilingual S2ST. Specifically, we decompose the S2ST task into speech-to-text translation (S2TT) and text-to-speech synthesis (TTS). For S2TT, we propose an effective speech language model that integrates the pretrained Whisper encoder for robust audio understanding and Qwen 3.0 for advanced text comprehension. A lightweight speech adapter is employed to bridge the modality gap between speech and text representations. To further facilitate the multimodal knowledge learning, a two-stage fine-tuning strategy is introduced. In the TTS stage, we adopt a streaming autoregressive generation approach to produce natural and fluent target speech. Experiments on the CVSS benchmark show that S2ST-Omni consistently outperforms existing state-of-the-art S2ST systems in translation quality, highlighting its effectiveness and superiority.
Index Terms—  Speech-to-speech translation, speech-text alignment, speech language model, streaming speech generation
$\dagger$ denotes the corresponding author.
1 Introduction
Multilingual speech-to-speech translation (S2ST) aims to directly convert utterance from multiple source languages into a target language speech, which plays a pivotal role in cross-language communication, education, and medical care [1, 2].
Conventional S2ST systems [3, 4, 5] normally adopt a statistically based cascaded architecture that decomposes the S2ST task into separate modules: automatic speech recognition (ASR) [6, 7], machine translation [8, 9], and text-to-speech synthesis (TTS) [10, 11]. Notable examples include the Verbmobil project [5], which employs a hybrid multi-engine approach to handle multilingual business dialogues, and the corpus-based framework proposed by [4], integrating multilingual speech recognition, text translation, and speech synthesis to achieve translation among English, Japanese, and Chinese. Nonetheless, cascaded pipelines inherently suffer from error propagation, rigid module interfaces, and limited real-time performance, restricting their effectiveness in complex practical scenarios [12].
With recent advances in deep learning, direct S2ST methods have emerged [13, 14, 12, 15], demonstrating better performance compared to previous methods. TranSpeech [13] introduced a non-autoregressive (NAR) speech-to-speech translation model that addresses acoustic multimodality through a bilateral perturbation mechanism that combines style normalization and information enhancement to learn deterministic linguistic representations. Afterwards, DASpeech [14] used a two-pass NAR architecture that first generates target text using a directed acyclic graph-based linguistic decoder and then produces target speech with a FastSpeech 2 [16] acoustic decoder, showcasing the effectiveness and potential of unified models to capture cross-lingual speech mappings. Despite impressive performance, these methods still rely heavily on large-scale aligned parallel speech corpora, limiting scalability to low-resource languages. Moreover, there remains substantial room to improve the translation accuracy. To mitigate this limitation, ComSpeech [12] introduced a compositional framework that leverages existing S2TT and TTS models, reducing dependence on triplet data. However, due to its reliance on a vocabulary adaptor and phoneme-level embeddings, ComSpeech generates relatively shallow phonetic representations with limited contextual integration. This inherent limitation reduces its capacity to model long-range dependencies, thereby constraining its effectiveness in handling semantically complex or ambiguous utterances.
Fig. 1: Overall architecture of the proposed S2ST-Omni framework. LSA denotes the lightweight speech adapter.
To solve these problems, we propose S2ST-Omni, an innovative and scalable framework for multilingual S2ST. Essentially, the core principle of S2ST-Omni lies in the efficient and seamless integration of large-scale multimodal pre-trained models, which can significantly reduce reliance on parallel speech data while enabling efficient and high-performance S2ST. To elaborate, we first present a speech language model (SLM) that leverages powerful pre-trained Whisper-large-v3 [17] encoder for robust audio comprehension and Qwen3.0 4B large language model (LLM) [18] for advanced text understanding. Meanwhile, a lightweight speech adapter (LSA) module is introduced to bridge the modality gap between speech and text, enabling efficient fusion and joint modeling of multimodal knowledge. To further enhance multimodal learning, a two-stage fine-tuning strategy is introduced. In the first stage, only the LSA module is updated to achieve initial convergence. Afterwards, both the LSA and Qwen LLM are jointly optimized during the second phase, where the LLM is fine-tuned using the parameter-efficient low-rank adaptation (LoRA) [19] method. Moreover, to improve translation quality and ensure low-latency speech synthesis, we adopt a pre-trained streaming speech generation approach [20] to synthesize the target utterance in an autoregressive (AR) fashion. Extensive experiments on the CVSS [21] corpus demonstrate that S2ST-Omni achieves substantial improvements over current state-of-the-art (SOTA) baselines, consistently delivering the best translation quality.
2 METHODOLOGY
In this section, we delineate the detailed architecture of the proposed S2ST-Omni framework, whose overall schematic is illustrated in Fig. 1.
2.1 System Overview
As shown in Fig. 1, our proposed S2ST-Omni framework comprises four key components: a pre-trained speech encoder, a learnable LSA, a pre-trained LLM, and a pre-trained streaming speech generation model. Specifically, to achieve accurate S2TT, we first leverage the pre-trained Whisper model in conjunction with a learnable LSA module to effectively extract and align speech representations. These features are then transformed into representations compatible with the pre-trained Qwen3.0 LLM. The entire S2TT model is optimized using cross-entropy loss to ensure accurate and efficient speech understanding and recognition. For efficient and fast speech synthesis, we employ a pre-trained streaming speech generation model [20] that generates target audio in an AR manner.
2.2 SLM via Seamless Speech-Text Alignment
Benefiting from training on large-scale and weakly supervised multilingual corpora, the Whisper [17] architecture has shown a strong generalization ability in capturing diverse acoustic and linguistic characteristics across a wide range of languages and conditions, including low-resource scenarios. These properties make it an ideal choice for multilingual speech understanding and translation tasks. Consequently, we adopt Whisper-large-v3 as the speech encoder in the proposed SLM. Let $X_{s}\in\mathbb{R}^{T}$ denotes the input source utterance, where $T$ is the duration of $X_{S}$, $\bm{T}\in\mathbb{R}^{L}$ represent its corresponding target transcript, where $L$ equals its text token length. The Whisper model first converts the raw waveform into a log-Mel spectrogram feature $F\in\mathbb{R}^{T_{1}\times D}$, where $T_{1}$ is the number of spectrogram frames and $D$ is the acoustic feature dimension. Then, the log-Mel spectrogram $F$ is passed through the Whisper encoder, which consists of a stack of transformer blocks [22] comprising multi-head self-attention and feed-forward layers. The encoder also incorporates positional embeddings to retain temporal information and layer normalization for stable training. Through this architecture, we can extract high-level speech representations that capture both local acoustic features and long-range contextual dependencies.
Subsequently, to fully exploit the powerful semantic understanding capabilities of the LLM, it is essential to bridge the representation gap between the Whisper encoder and the Qwen 3.0 LLM. To this end, we present a LSA module $\mathcal{A}(\cdot)$ based on a multilayer projection (MLP) module, which maps the encoded speech features into a compatible representation space compatible with the Qwen 3.0 LLM. Formally, assume $\mathcal{F}_{\text{whisper}}\in\mathbb{R}^{T_{w}\times D_{w}}$ denote the encoded speech features produced by the Whisper encoder. These features are further transformed by a four-layer feed-forward speech adapter $\mathcal{A}(\cdot)$, yielding the adapted representation $\mathcal{F}_{\text{adapt}}$ as:
$\begin{split}\mathcal{H}_{0}&=\mathcal{F}_{\text{whisper}}\end{split}$ (1)
$\mathcal{H}_{i}=\sigma(\mathcal{H}_{i-1}\mathbf{W}_{i}+\mathbf{b}_{i}),\quad\forall i\in\{1,2,3,4\}$ (2)
$\begin{split}\mathcal{F}_{\text{adapt}}&=\mathcal{H}_{4},\end{split}$ (3)
where $\mathcal{F}_{\text{adapt}}\in\mathbb{R}^{T_{a}\times D_{a}}$, $\mathbf{W}_{i}$ and $\mathbf{b}_{i}$ are learnable weight and bias parameters of the $i$-th linear layer, $\sigma(\cdot)$ denotes the LeakyReLU activation function with a negative slope of 0.1, and the operator $\prod$ represents a nested functional composition. In this way, S2ST-Omni enables efficient alignment of speech features with the language model’s input space while maintaining computational simplicity and training stability.
Finally, the adjusted representation $\mathcal{F}_{\text{adapt}}$ is input into the Qwen 3.0 LLM for S2TT, under the guidance of the conventional cross-entropy loss:
$\mathcal{L}_{\text{SLM}}=-\sum_{l=1}^{L}\log P(t_{l}\mid t_{<l},\mathcal{F}_{\text{adapt}})$ (4)
where ${t_{l}}$ is the $l$-th predicted target text token, and $t_{<l}$ is the autoregressively generated text tokens preceding the $l$-th position.
2.3 Two-Stage Fine-tuning Strategy
To further improve the translation performance of the proposed S2ST-Omni, we introduce an effective two-stage progressive fine-tuning strategy, which is illustrated in Fig. 1.
In the first stage, we freeze the parameters of the pretrained Whisper encoder and the Qwen 3.0 LLM, and update merely the LSA module. Generally, this phase focuses on learning an effective alignment between the acoustic representations produced by Whisper and the textual embeddings of the LLM, without disrupting the pretrained knowledge encoded in the multimodal backbone components. Once S2ST-Omni reaches convergence in the first phase, we jointly optimize both the LSA and the Qwen LLM, where the LLM is fine-tuned using the parameter-efficient low-rank adaptation (Lora) mechanism. Concretely, LoRA is applied to the key projection matrices, i.e., q_proj and v_proj within the attention module, effectively reducing training overhead while maintaining performance. By this means, it enables the LLM to better adapt to speech-derived representations and enhances its ability to handle complex linguistic phenomena in spoken language, such as disfluencies, code-switching, and semantic ambiguities.
2.4 Streaming Speech Generation
To achieve a balance between synthesis latency and translation quality, we incorporate a pretrained streaming speech generation method, CosyVoice2 [20]. To be specific, we first use a supervised speech tokenizer1 to convert the target waveform $X_{T}$ into a sequence of discrete speech tokens $Y=[y_{1},y_{2},\ldots,y_{M}]$, sampled at a rate of 25 tokens per second, where each token $y_{i}\in\{0,1,\ldots,6561\}$. The predicted text output from the Qwen 3.0 LLM, denoted as $\bm{\hat{T}}$, is subsequently fed into a text-to-speech language model, which autoregressively generates the predicted speech token sequence $\hat{\mathbf{Y}}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{M}]$. Following this, $\hat{\mathbf{Y}}$ is reconstructed into target Mel-spectrogram features $\bm{\hat{\mathbf{F}}}_{\mathrm{mel}}$ using a chunk-aware causal conditional flow matching (CFM) model [23, 24, 25], conditioned on prompt speaker embeddings. In particular, every $K$ generated speech tokens are grouped as a chunk, which serves as the input unit for Mel-spectrogram generation. Last, the target Mel–spectrograms are fed into a pretrained HifiGAN Vocoder2 to generate the converted utterance. In our case, we directly utilize the pretrained CFM3 and HifiGAN without any fine-tuning.
1https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/speech_tokenizer_v2.onnx2https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/flow.pt3https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B/blob/main/hift.pt
3 EXPERIMENTS
In this section, we first describe the datasets used for the S2ST task. We then present the implementation details of the proposed S2ST-Omni model and briefly introduce the baseline systems used for comparison. Finally, we report and analyze the overall experimental results.
3.1 Experimental Settings
3.1.1 Databases
[…]
Table 1: ASR-BLEU score comparison among all evaluated S2ST methods.
[…]
S2UT [26] 22.23 2.99 18.53 14.58
Translatotron 2 [27] 26.12 16.92 22.98 22.01
DASpeech [14] 25.03 16.14 21.37 20.85
UnitY [28] 26.90 16.36 24.06 22.44
ComSpeech [12] 28.15 18.16 24.80 23.70
StreamSpeech [15] 28.45 20.93 27.25 25.54
S2ST-Omni 30.20 32.09 28.02 30.10
3.1.2 Implementation Details
For all experiments, the proposed S2ST-Omni system is trained using the AdamW optimizer with an initial learning rate of $1\times 10^{-5}$ and a batch size of 32. Training is conducted for 100,000 steps in each of the two fine-tuning stages within the PyTorch framework, utilizing 8 NVIDIA A100 GPUs. The system prompt for the Qwen 3.0 4B LLM is defined as: “You are an expert in speech-to-speech translation, responsible for accurately translating multilingual input speech into corresponding English text.”
Regarding the evaluating metric, we employ the commonly used ASR-BLEU to assess the performance of all comparative approaches. In practice, the translated speech is first transcribed into its corresponding transcript using a pre-trained ASR model4, after which the BLEU score [29] is computed using the SacreBLEU toolkit5 to assess translation quality and statistical significance.
4https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_960h_pl.pt5https://github.com/mjpost/sacrebleu
3.1.3 Baselines
To ensure a fair and comprehensive comparison, we evaluate our S2ST-Omni system against six well-established S2ST methods:
  • S2UT [26]: A single-stage NAR model that directly translates speech into discrete units.
  • Translatotron 2 [27]: An end-to-end, one-pass system for direct speech-to-speech translation.
  • DASpeech [14]: A two-stage non-autoregressive framework that first converts speech to text and subsequently from text to discrete units.
  • UnitY [28]: A fully AR pipeline employing two passes to improve translation accuracy.
  • ComSpeech [12]: A composite model integrating pretrained S2TT and TTS modules into a direct S2ST system via contrastive learning.
  • StreamSpeech [15]: A unified multi-task model that performs direct simultaneous S2ST by jointly learning translation and timing policies for real-time streaming input.
3.2 Results and Analysis
To validate the effectiveness of the proposed approach, we compare the translation performance of S2ST-Omni against six state-of-the-art S2ST systems. A comprehensive summary of the results is provided in Table 1.
As shown in Table 1, S2ST-Omni consistently and significantly outperforms all baselines on the CVSS benchmark across the evaluated language pairs, with over 50% relative ASR-BLEU improvements in certain cases. These results highlight the strong generalization capability of the proposed framework across linguistically diverse settings. To be concrete, in the French $\rightarrow$ English direction, our S2ST-Omni achieves an ASR-BLEU score of 30.20, surpassing the strongest baseline (StreamSpeech, 28.45). This substantial improvement underscores the robustness of the propose method in handling phonetic variability and prosodic features characteristic of French. For German $\rightarrow$ English, the S2ST-Omni attains a 28.02 ASR-BLEU score, the highest among all language pairs as well, demonstrating its effectiveness in managing syntactic complexity and compound structures inherent to German. In the Spanish $\rightarrow$ English setting, S2ST-Omni achieves a remarkable 32.09 ASR-BLEU score, again outperforming all baselines by a wide margin. This result reflects the model’s ability to capture the rich morphological patterns of Spanish and generate fluent, accurate translations. In addition, despite the significant differences in phonology, syntax, and morphology among French, Spanish, and German, S2ST-Omni consistently delivers strong and comparable ASR-BLEU scores, further validating its stability and robustness across diverse linguistic scenarios.
4 CONCLUSIONS
In this study, we introduce an efficient and scalable framework, S2ST-Omni, for multilingual speech-to-speech translation. First, we propose an effective speech language model that integrates large-scale pre-trained multimodal models (Whisper and Qwen 3.0) with a lightweight speech adapter to achieve high-quality translation while reducing reliance on parallel speech data. Then, a two-stage progressive fine-tuning approach is presented to further improve the translation accuracy. Additionally, we employ a streaming speech generation model to balance the trade-off between the synthesis latency and translation quality. Extensive experiments on the CVSS benchmark confirm that S2ST-Omni achieves better translation quality than existing state-of-the-art systems, demonstrating its effectiveness and superiority.